# 2D-OOB: Attributing Data Contribution through Joint Valuation Framework

### v1

@@ Abstract
Data valuation has emerged as a powerful framework to quantify the contribution of each datum to the training of a particular machine learning model. However, it is crucial to recognize that the quality of various cells within a single data point can vary greatly in practice. For example, even in the case of an abnormal data point, not all cells are necessarily noisy. The single scalar valuation assigned by existing methods blurs the distinction between noisy and clean cells of a data point, thereby compromising the interpretability of the valuation. In this paper, we propose 2D-OOB, an out-of-bag estimation framework for jointly determining helpful (or detrimental) samples, as well as the particular cells that drive them. Our comprehensive experiments demonstrate that 2D-OOB achieves state-of-the-art performance across multiple use cases, while being exponentially faster. 2D-OOB excels in detecting and rectifying fine-grained outliers at the cell level, as well as localizing backdoor triggers in data poisoning attacks.
@@ 1 Introduction
From customer behavior prediction and medical image analysis to autonomous driving and policy making, machine learning (ML) systems process ever increasing amounts of data. In such data-rich regimes, a fraction of the samples is often noisy, incorrect annotations are likely to occur, and uniform data quality standards become difficult to enforce. To address these challenges, data valuation emerges as a research field receiving increasing attention, focusing on properly assessing the contribution of each datum to ML training [12]. These methods have proven useful in identifying low-quality samples that can be detrimental to model performance, as well as selecting subsets of data that are representative of enhanced model performance [23, 48, 27]. Furthermore, they are widely applicable in data marketplace for fair revenue allocation and incentive design [51, 45, 40].
Nevertheless, existing data valuation methods assign a scalar score to each datum, thereby failing to account for the varied roles of individual cells. This leaves the valuation rationale unclear and can be unsatisfactory and sub-optimal in various practical scenarios. Firstly, whenever a score is assigned to a data point by a particular data valuation method, it is crucial to understand the underlying justifications to ensure transparency and reliability, especially in high-stakes decision making [39]. Secondly, it is important to recognize the fact that even if a data point is of low quality, it is rarely the case that all the cells within this data point are noisy [37, 26, 43]. The absence of detailed insights into how individual cells contribute to ML training inevitably leads to discarding entire data points. This can result in substantial data waste, particularly when only a few cells are noisy and data acquisition is expensive. Finally, in data markets, different cells within a data point may originate from different data sellers [3, 10]. Consequently, a singular valuation for the entire point fails to offer equitable compensation to all contributing parties.
Figure 1: Comparison of data valuation and joint valuation. (a) Data valuation evaluates the quality of individual data points, whereas (b) joint valuation evaluates the quality of individual cells. Both panels illustrate the same hypothetical dataset, and the darker colors overlaid represent the higher quality or importance. Joint valuation provides a finer level of attributions than data valuation and aims to describe how features affect data values. As panel (b) illustrates, the joint valuation framework can identify outlier cells highlighted with blue boxes (i.e., $-1$ in "Income" and $100$ in "Education") and provide quantitative interpretations of data values.
@@
In this paper, we propose 2D-OOB, a powerful and efficient joint valuation framework that can attribute a data point’s value to its individual features. 2D-OOB quantifies the importance of each cell in a dataset, as illustrated in Figure 1, providing interpretable insights into which cells are associated with influential data points. Our method is computationally efficient as well as theoretically supported by its connections with Data-OOB [23]. Moreover, our extensive empirical experiments demonstrate the practical effectiveness of 2D-OOB in various use cases. 2D-OOB accurately identifies cell outliers and pinpoints which cells to fix to improve model performance. 2D-OOB enables inspection of data poisoning attacks by precisely localizing the backdoor trigger, an artifact inserted into a training sample to induce malicious model behavior [13, 5]. 2D-OOB is on average $200$ times faster than state-of-the-art methods across all datasets examined.
@@ 2 Preliminaries
The primary goal of data valuation is to quantify the contribution of individual data points to a model’s performance. Leveraging the Shapley value in cooperative game theory [38], DataShapley [12] measures the average change in a utility function $U:2^{\mathcal{D}}\rightarrow\mathbb{R}$ when a data point is removed. For $i\in[n]$, DataShapley of $i$-th datum is defined as follows.
$\phi_{i}^{\mathrm{Shap}}:=\frac{1}{n}\sum_{k=1}^{n}\frac{1}{\binom{n-1}{k-1}}\sum_{S\subset\mathcal{D}_{k}^{(i)}}[U(S\cup\{(x_{i},y_{i})\})-U(S)]$ (1)
where $\mathcal{D}_{k}^{(i)}:=\{S\subseteq\mathcal{D}|(x_{i},y_{i})\notin S,|S|=k-1\}$. DataShapley $\phi_{i}^{\mathrm{Shap}}$ in (1) considers every set $S\in\mathcal{D}_{k}^{(i)}$ and computes the average difference in utility $U(S\cup\{(x_{i},y_{i})\})-U(S)$. It characterizes the impact of a data point, but its computation requires evaluating $U$ for all possible subsets of $\mathcal{D}$, rendering precise calculations infeasible. Many efficient computation algorithms have been studied [15, 25, 50], and in these studies, Shapley-based methods have demonstrated better effectiveness in detecting low-quality samples than standard attribution approaches, such as leave-one-out and influence function methods [20, 9].
@@
As an alternative efficient data valuation method, Kwon and Zou [23] propose Data-OOB, which leverages a bagging model and measures the similarity between a nominal label and weak learners’ predictions. To be more specific, we suppose a bagging model consists of $B$ weak learners, where for $b\in[B]$, the $b$-th weak learner $\hat{h}_{b}$ is given as a minimizer of the weighted empirical risk,
@@
It is noteworthy that Data-OOB in (2) can be computed by training a single bagging model, making it computationally efficient. Kwon and Zou [23] show that Data-OOB can easily scale to millions of data points, but for DataShapley this is often very impractical. In addition, Data-OOB is often comparable to or even more effective than DataShapley in detecting mislabeled data points and selecting helpful data points [23, 16].
@@ 3 Attributing Data Contribution through Joint Valuation Framework
Data valuation quantifies desiderata of data points, however, it does not describe what features contribute and how much to those specific data values. For instance, in anomaly detection tasks, data valuation methods can be deployed to detect anomalous data points, but they do not explain why they are abnormal, which is not generally desirable in practice. To address this challenge, we consider a joint valuation framework and assess a cell score for each feature of a data point. Here, a cell score is designed to quantify how a feature affects the value of an individual data point, attributing a data value to features.
To the best of the author’s knowledge, Liu et al. [29] first consider a concept of the joint valuation in literature and introduce 2D-Shapley to quantitatively interpret DataShapley. To this end, we denote a 2D utility function by $u:[n]\times[d]\to\mathbb{R}$, which takes as input a subset of data points $S\subseteq[n]$ and a subset of features $F\subseteq[d]$, and measure the utility of a fragment of the given dataset consisting of cells $\{(i,j)\}_{i\in S,j\in F}$, where a tuple $(i,j)$ denotes a cell at the $i$-th datum and the $j$-th column. Then, 2D-Shapley is defined as
$\psi_{ij}^{\mathrm{2D-Shap}}:=\frac{1}{nd}\sum_{k=1}^{n}\sum_{l=1}^{d}\frac{1}{\binom{n-1}{k-1}\binom{d-1}{l-1}}\sum_{(S,F)\subset\mathcal{D}_{k,l}^{(i,j)}}M_{u}^{i,j}(S,F)$ (3)
@@
The function $M_{u}^{i,j}$ allows us to quantify how much removing a specific cell at $(i,j)$ from a given set $(S\cup\{i\},F\cup\{j\})$ affects the overall utility, and 2D-Shapley $\psi_{ij}^{\mathrm{2D-Shap}}$ evaluates the average $M_{u}^{i,j}$ across all possible data fragments $(S,F)\subset\mathcal{D}_{k,l}^{(i,j)}$.
@@
We propose 2D-OOB, an efficient and model-agnostic joint valuation framework that leverages out-of-bag estimation to attribute data contribution. We further illustrate how 2D-OOB is connected to Data-OOB, thereby facilitating sample-wise interpretation for data valuation in Section 3.2.
@@
where $x_{i,S_{b}}$ is a subvector of $x_{i}$ that only takes elements in a subset $S_{b}$. This difference enables us to assess the impact of which features are more influential: if $S_{b}$ includes a helpful (or detrimental) feature, we can expect the out-of-bag prediction $\hat{f}(x_{i,S_{b}})$ to be good (or poor). We formalize this intuition and propose 2D-OOB. For $i\in[n]$, $j\in[d]$ and $\{(w_{b},S_{b},\hat{f}_{b})\}_{b=1}^{B}$, the 2D-OOB for the $j$-th cell of the $i$-th data point is defined as follows,
@@
where $T:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ is a utility function that scores the performance of the weak learner $\hat{f}_{b}(x_{i,S_{b}})$ on the $i$-th datum $(x_{i},y_{i})$. Specifically, for binary or multi-class classification problems, we can adopt $T(y_{i},\hat{f}_{b}(x_{i,S_{b}}))=\mathds{1}(y_{i}=\hat{f}_{b}(x_{i,S_{b}}))$. In this case, 2D-OOB measures the average accuracy score of out-of-bag predictions (specifically, when the $i$-th data point is out-of-bag) if a cell $j$ is used in training $\hat{f}_{b}$. For regression problems, we can use the negative squared error loss function, defined as $T(y_{i},\hat{f}_{b}(x_{i,S_{b}}))=-(y_{i}-\hat{f}_{b}(x_{i,S_{b}}))^{2}$. In practice, $\mathcal{X}$ could also be incorporated into $T$ to suit the specific use case.
While Data-OOB in (2) aims to assess the impact of the $i$-th datum, 2D-OOB in (4) provides interpretable insights by evaluating the data point with various combinations of features, revealing which cells are influential to model performance. Leveraging subset bagging scheme, 2D-OOB requires a single training of the bagging model, and thus it is computational efficiency.
@@ 3.2 Connection to Data-OOB
A proof is given in the Appendix C. Proposition 3.1 shows that 2D-OOB $\psi_{ij}^{\mathrm{2D-OOB}}$ can be expressed as a conditional empirical expectation of Data-OOB provided that the $j$-th feature is used in Data-OOB computation. It provides intuitive interpretations: for a fixed $i$ and $j\neq k$, $\psi_{ij}^{\mathrm{2D-OOB}}>\psi_{ik}^{\mathrm{2D-OOB}}$ implies that the cell $x_{ij}$ is more helpful to achieve the high OOB score, which serves as an indicator of model performance, than the cell $x_{ik}$.
@@ 4 Experiments
In this section, we empirically show the effectiveness of 2D-OOB across multiple use cases of the joint valuation: cell-level outlier detection, cell fixation, and backdoor trigger detection. As a summary, 2D-OOB can precisely identify anomalous cells that should be prioritized for examination and subsequent fixation to improve model performance. In the context of backdoor trigger detection, 2D-OOB demonstrates its efficacy by accurately identifying different types of triggers within poisoned data, showcasing its proficiency in detecting non-random, targeted anomalies. Our method also exhibits high computational efficiency through run-time comparison.
Throughout all of our experiments, 2D-OOB uses a subset bagging model with $B=1000$ decision trees. We randomly select a fixed ratio of features to build each decision tree. Unless otherwise specified, we utilize half of the features for each weak learner and set $T(y_{i},\hat{f}(x_{i,S_{b}}))=\mathds{1}(y_{i}=\hat{f}(x_{i,S_{b}}))$. The run time is measured on a single Intel Xeon Gold 6226 2.9 Ghz CPU processor.
@@ Results
We also evaluate the area under the curve (AUC) as a quantitative metric and the run-time. As Table 1 shows, 2D-OOB achieves an average AUC of $0.83$ across $12$ datasets, compared to $0.67$ for 2D-KNN, while being significantly faster. For high-dimensional datasets such as the musk dataset, which comprises $166$ features, 2D-KNN would take more than an hour to process, while 2D-OOB can finish in seconds. Furthermore, we present additional results on multi-class classification datasets in Appendix B.1, demonstrating the consistently superior performance and efficiency of 2D-OOB.
Table 1: Cell-level outlier detection results. AUC and run-time comparison between 2D-OOB and 2D-KNN across the twelve datasets. The average and standard error of the AUC and run-time (in seconds) based on $30$ independent experiments are denoted by “average $\pm$ standard error”. Bold numbers denote the best method. The AUC value for the Random method consistently remains at $0.5$ across all datasets. Overall, 2D-OOB achieves a significantly higher AUC while being orders of magnitude faster than 2D-KNN.

### v2

@@ Abstract
Data valuation has emerged as a powerful framework for quantifying each datum’s contribution to the training of a machine learning model. However, it is crucial to recognize that the quality of cells within a single data point can vary greatly in practice. For example, even in the case of an abnormal data point, not all cells are necessarily noisy. The single scalar score assigned by existing data valuation methods blurs the distinction between noisy and clean cells of a data point, making it challenging to interpret the data values. In this paper, we propose 2D-OOB, an out-of-bag estimation framework for jointly determining helpful (or detrimental) samples as well as the particular cells that drive them. Our comprehensive experiments demonstrate that 2D-OOB achieves state-of-the-art performance across multiple use cases while being exponentially faster. Specifically, 2D-OOB shows promising results in detecting and rectifying fine-grained outliers at the cell level, and localizing backdoor triggers in data poisoning attacks.
@@ 1 Introduction
From customer behavior prediction and medical image analysis to autonomous driving and policy making, machine learning (ML) systems process ever increasing amounts of data. In such data-rich regimes, a fraction of the samples is often noisy, incorrect annotations are likely to occur, and uniform data quality standards become difficult to enforce. To address these challenges, data valuation emerges as a research field receiving increasing attention, focusing on properly assessing the contribution of each datum to ML training [12, 24]. These methods have proven useful in identifying low-quality samples that can be detrimental to model performance, as well as selecting subsets of data that are representative of enhanced model performance [50, 28, 47]. Furthermore, they are widely applicable in data marketplace for fair revenue allocation and incentive design [53, 46, 41].
Nevertheless, existing data valuation methods assign a scalar score to each datum, thereby failing to account for the varied roles of individual cells. This leaves the valuation rationale unclear and can be unsatisfactory and sub-optimal in various practical scenarios. Firstly, whenever a score is assigned to a data point by a particular data valuation method, it is crucial to understand the underlying justifications to ensure transparency and reliability, especially in high-stakes decision making [40]. Secondly, it is important to recognize the fact that even if a data point is of low quality, it is rarely the case that all the cells within this data point are noisy [38, 27, 44]. The absence of detailed insights into how individual cells contribute to ML training inevitably leads to discarding the entire data point. This can result in substantial data waste, particularly when only a few cells are noisy and data acquisition is expensive. Finally, in data markets, different cells within a data point may originate from different data sellers [3, 10]. Consequently, a singular valuation for the entire point fails to offer equitable compensation to all contributing parties.
Figure 1: Comparison of data valuation and joint valuation. (a) Data valuation evaluates the quality of individual data points, whereas (b) joint valuation evaluates the quality of individual cells. Both panels illustrate the same hypothetical dataset, while darker colors indicate higher quality or importance. As illustrated in panel (a), data valuation can only identify that the third and fifth data points are of low quality, but it lacks further feature-level attribution. This limitation may result in discarding the entire data point, even when only certain cells are problematic. In contrast, joint valuation provides a finer level of attribution than data valuation and aims to reveal how individual features contribute to data values. As shown in panel (b), the joint valuation framework can identify outlier cells (highlighted by blue boxes), such as $-1$ in "Income" and $100$ in "Education", providing detailed interpretations of data values.
@@
In this paper, we propose 2D-OOB, a powerful and efficient joint valuation framework that can attribute a data point’s value to its features. 2D-OOB quantifies the importance of each cell in a dataset, as illustrated in Figure 1, providing interpretable insights into which cells are associated with influential data points. Our method is computationally efficient as well as theoretically supported by its connections with Data-OOB [24]. Moreover, our extensive empirical experiments demonstrate the practical effectiveness of 2D-OOB in various use cases. 2D-OOB accurately identifies cell outliers and pinpoints which cells to fix to improve model performance. Additionally, 2D-OOB enables inspection of data poisoning attacks by precisely localizing the backdoor trigger, an artifact inserted into a training sample to induce malicious model behavior [13, 5]. On average, 2D-OOB is $200$ times faster than state-of-the-art methods across all datasets examined.
@@ 2 Preliminaries
The primary goal of data valuation is to quantify the contribution of individual data points to a model’s performance. Leveraging the Shapley value in cooperative game theory [39], DataShapley [12] measures the average change in a utility function $U:2^{\mathcal{D}}\rightarrow\mathbb{R}$ over all possible subsets of the dataset that either include or exclude the data point. For $i\in[n]$, DataShapley of $i$-th datum is defined as follows.
$\phi_{i}^{\mathrm{Shap}}:=\frac{1}{n}\sum_{k=1}^{n}\frac{1}{\binom{n-1}{k-1}}\sum_{S\in\mathcal{D}_{k}^{(i)}}[U(S\cup\{(x_{i},y_{i})\})-U(S)],$ (1)
where $\mathcal{D}_{k}^{(i)}:=\{S\subseteq\mathcal{D}|(x_{i},y_{i})\notin S,|S|=k-1\}$. DataShapley $\phi_{i}^{\mathrm{Shap}}$ in (1) considers every set $S\in\mathcal{D}_{k}^{(i)}$ and computes the average difference in utility $U(S\cup\{(x_{i},y_{i})\})-U(S)$. It characterizes the impact of a data point, but its computation requires evaluating $U$ for all possible subsets of $\mathcal{D}$, rendering precise calculations infeasible. Many efficient computation algorithms have been proposed [15, 26, 52], and in these studies, Shapley-based methods have demonstrated better effectiveness in detecting low-quality samples than standard attribution approaches, such as leave-one-out and influence function methods [20, 9].
@@
As an alternative efficient data valuation method, Kwon and Zou [24] propose Data-OOB, which leverages a bagging model and measures the similarity between the nominal label and weak learners’ predictions. To be more specific, consider a bagging model consisting of $B$ weak learners, where for $b\in[B]$, the $b$-th weak learner $\hat{h}_{b}$ is given as a minimizer of the weighted empirical risk,
@@
It is noteworthy that Data-OOB in (2) can be computed by training a single bagging model, making it computationally efficient. Kwon and Zou [24] show that Data-OOB can easily scale to millions of data points, but for DataShapley this is often very impractical. In addition, Data-OOB is typically comparable to or even more effective than DataShapley in detecting mislabeled data points and selecting helpful data points [24, 16].
@@ 3 Attributing Data Contribution through Joint Valuation Framework
Data valuation quantifies the utility of data points, however, it fails to identify which features contribute to these data values and to what extent. For instance, in anomaly detection tasks, data valuation methods can be deployed to detect anomalous data points but cannot explain why they are considered abnormal, which is generally not desirable in practice. To address this challenge, we introduce a joint valuation framework that assigns a cell score to each feature of a data point. Here, a cell score is designed to quantify how each feature affects the value of an individual data point, thereby attributing the data value to specific features.
To the best of the author’s knowledge, Liu et al. [30] were the first to consider the concept of joint valuation in the literature, proposing 2D-Shapley as a means to quantitatively interpret DataShapley. To formalize this, we denote a 2D utility function by $u:[n]\times[d]\to\mathbb{R}$, which takes as input a subset of data points $S\subseteq[n]$ and a subset of features $F\subseteq[d]$, measuring the utility of a fragment of the given dataset consisting of cells $\{(i,j)\}_{i\in S,j\in F}$, where a tuple $(i,j)$ denotes a cell at the $i$-th datum and the $j$-th column. Then, 2D-Shapley is defined as
$\psi_{ij}^{\mathrm{2D-Shap}}:=\frac{1}{nd}\sum_{k=1}^{n}\sum_{l=1}^{d}\frac{1}{\binom{n-1}{k-1}\binom{d-1}{l-1}}\sum_{(S,F)\in\mathcal{D}_{k,l}^{(i,j)}}M_{u}^{i,j}(S,F)$ (3)
@@
The function $M_{u}^{i,j}$ allows us to quantify how much removing a specific cell at $(i,j)$ from a given set $(S\cup\{i\},F\cup\{j\})$ affects the overall utility, and 2D-Shapley $\psi_{ij}^{\mathrm{2D-Shap}}$ evaluates the average $M_{u}^{i,j}$ across all possible data fragments $(S,F)\in\mathcal{D}_{k,l}^{(i,j)}$.
@@
To address these limitations, we propose 2D-OOB, an efficient and model-agnostic joint valuation framework that leverages out-of-bag estimation to attribute data contribution. We also illustrate how 2D-OOB is connected to Data-OOB, thereby facilitating sample-wise feature-level interpretation for data valuation, as discussed in Section 3.2.
@@
where $x_{i,S_{b}}$ is a subvector of $x_{i}$ that only takes elements in a subset $S_{b}$. This difference enables us to assess the impact of which features are more influential: if $S_{b}$ includes a helpful (or detrimental) feature, we can expect the out-of-bag prediction $\hat{f}(x_{i,S_{b}})$ to be good (or poor). We formalize this intuition and propose 2D-OOB. For $i\in[n]$, $j\in[d]$ and $\{(\mathbf{w}_{b},S_{b},\hat{f}_{b})\}_{b=1}^{B}$, the 2D-OOB for the $j$-th cell of the $i$-th data point is defined as follows,
@@
where $T:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ is a utility function that scores the performance of the weak learner $\hat{f}_{b}(x_{i,S_{b}})$ on the $i$-th datum $(x_{i},y_{i})$. Specifically, for binary or multi-class classification problems, we can adopt $T(y_{i},\hat{f}_{b}(x_{i,S_{b}}))=\mathds{1}(y_{i}=\hat{f}_{b}(x_{i,S_{b}}))$. In this case, 2D-OOB measures the average accuracy score of out-of-bag predictions (specifically, when the $i$-th data point is out-of-bag) if the $j$-th feature is used in training $\hat{f}_{b}$. For regression problems, we can use the negative squared error loss function, defined as $T(y_{i},\hat{f}_{b}(x_{i,S_{b}}))=-(y_{i}-\hat{f}_{b}(x_{i,S_{b}}))^{2}$. In practice, $\mathcal{X}$ could also be incorporated into $T$ to suit the specific use case.
While Data-OOB in (2) aims to assess the impact of the $i$-th datum, 2D-OOB in (4) further provides interpretable insights by evaluating the data point with various combinations of features, revealing which cells are influential to model performance. By leveraging the subset bagging scheme, 2D-OOB only requires a single training of the bagging model, making it computationally efficient.
@@ 3.2 Connection to Data-OOB
A proof is given in Appendix C. Proposition 3.1 shows that 2D-OOB $\psi_{ij}^{\mathrm{2D-OOB}}$ can be expressed as a conditional empirical expectation of Data-OOB provided that the $j$-th feature is used in Data-OOB computation. It provides intuitive interpretations: for a fixed $i$ and $j\neq k$, $\psi_{ij}^{\mathrm{2D-OOB}}>\psi_{ik}^{\mathrm{2D-OOB}}$ implies that the cell $x_{ij}$ is more helpful to achieve a higher OOB score than the cell $x_{ik}$, where the OOB score serves as an indicator of model performance. By distinguishing the contributions of individual cells, 2D-OOB effectively realizes joint valuation, providing a finer granularity of analysis that links feature-level importance to individual data quality.
@@ 4 Experiments
In this section, we empirically show the effectiveness of 2D-OOB across multiple use cases of the joint valuation: cell-level outlier detection, cell fixation, and backdoor trigger detection. To the best of our knowledge, the latter two use cases are introduced here for the first time within the joint valuation framework. As a summary, 2D-OOB can precisely identify anomalous cells that should be prioritized for examination and subsequent fixation to improve model performance. In the context of backdoor trigger detection, 2D-OOB demonstrates its efficacy by accurately identifying different types of triggers within poisoned data, showcasing its proficiency in detecting non-random, targeted anomalies. Our method also exhibits high computational efficiency through run-time comparison.
Throughout all of our experiments, 2D-OOB uses a subset bagging model with $B=1000$ decision trees. We randomly select a fixed ratio of features to build each decision tree. Unless otherwise specified, we utilize half of the features for each weak learner and set $T(y_{i},\hat{f}(x_{i,S_{b}}))=\mathds{1}(y_{i}=\hat{f}(x_{i,S_{b}}))$. The run time is measured on a single Intel Xeon Gold 6226 2.9 GHz CPU processor. We provide a detailed ablation study of key hyperparameters in Section 4.4.
@@ Results
We also evaluate the area under the curve (AUC) as a quantitative metric and measure the run-time. As Table 1 shows, 2D-OOB achieves an average AUC of $0.83$ across $12$ datasets, compared to $0.67$ for 2D-KNN, while being significantly faster. For high-dimensional datasets such as the musk dataset, which comprises $166$ features, 2D-KNN would take more than an hour to process, while 2D-OOB can finish in seconds. Furthermore, we present additional results on multi-class classification datasets in Appendix B.1, demonstrating the consistently superior performance and efficiency of 2D-OOB.
Table 1: Cell-level outlier detection results. AUC and run-time comparison between 2D-OOB and 2D-KNN across twelve binary classification datasets. The average and standard error of the AUC and run-time (in seconds) based on $30$ independent experiments are denoted by “average $\pm$ standard error”. Bold numbers denote the best method. The AUC value for the Random method consistently remains at $0.5$ across all datasets. In every dataset, 2D-OOB achieves a significantly higher AUC while being orders of magnitude faster than 2D-KNN.